pooja195 (pooja195):

Glenna wants to rent a car for a trip to Key West for one week. She calls two car rental companies to get prices. Mr. Kotter’s Rentals rents a Cadillac Escalade for $99 for one week plus $0.11 per mile over 100 miles. Barbarino’s Rentals has the same vehicle for $75 per week and $0.15 per mile over 150 miles. If it is 432 miles to Key West, which company has the better deal? How many miles would Glenna drive before she would be spending the same amount at either company? @e.mccormick

OpenStudy (anonymous):

I'm too lazy to do it but poona ur smarter than most

OpenStudy (sleepyhead314):

(amount per week) + (amount per mile)*(432 - "over _#_")

OpenStudy (sleepyhead314):

you closed the question @pooja195 do you still need help?

pooja195 (pooja195):

yes lol i didnt want attention attracted to it i get confused when to may people answer but im totally lost on this

OpenStudy (sleepyhead314):

did you kinda get the format I gave above or would you like for me to rephrase it? :)

pooja195 (pooja195):

hmmm can you please rephrase it :)

OpenStudy (sleepyhead314):

you keep the flat number away from the variable so the flat number in this case is "$ per week" (since you are only renting for a week) like y = mx + b "$ per week" is b

pooja195 (pooja195):

y=.11/100+75?

OpenStudy (sleepyhead314):

you are going to try to make two different equations to compare them :P

OpenStudy (sleepyhead314):

m = rate which will be the "$ per mile"

OpenStudy (sleepyhead314):

x = (432 - "over ___ miles")

OpenStudy (sleepyhead314):

y = mx + b

pooja195 (pooja195):

100?

OpenStudy (sleepyhead314):

umm yes the "over ___ miles" for the first equation will be 100

pooja195 (pooja195):

ok so y= 100/1x-432?

OpenStudy (sleepyhead314):

hmmm "$ per week" = b "$ per mile" = m (432 - "over ___ miles") = x for y = mx + b for the first set of information b = "$ per week" = 99 m = "$ per mile" = 0.11 x = (432 - "over ___ miles") = (432 - 100) so y = (432 - 100)*(0.11) + 99

pooja195 (pooja195):

O_O

OpenStudy (anonymous):

:(

OpenStudy (sleepyhead314):

where'd I loose you, sorry

OpenStudy (sleepyhead314):

y = (432 - 100)*(0.11) + 99 = 135.52 btw

pooja195 (pooja195):

Ok i used that and the secound equation i had already :P forgot to put it and came up with Barbarino’s. Glenna would have to drive 887.5 miles to have both deals be the same.

OpenStudy (sleepyhead314):

no, not what I got...

OpenStudy (sleepyhead314):

unless I'm reading it wrong does the "$0.15 per mile over 150 miles." mean that only 0.15 will be added per mile *after* 150 miles have been driven?